# The NEMESIS Catalogue of Young Stellar Objects for the Orion Star Formation Complex

### v1

@@ 4.3.3 Gaia’s Renormalised Unit Weight Error
4.4 Big data identification of visual-pair candidates
In Sec. 2.4 we have employed the VizieR-SED tool to retrieve the photometric data available on CDS around the position of each source in our catalogue. In this section, we further employ these data to gauge sources’ neighbourhoods.
Starting from the raw VizieR-SED output for a 5” searching radius, rather than relying again on the procedure described in Sec. 3.2 we carried a simpler preprocessing. We kept only data in the wavelength range 0.01-5, excluded multi-epoch surveys (e.g., TESS) and surveys with source detection and extraction at different apertures (e.g., SDSS). We then grouped the remaining data by unique CDS table name and photometric band and counted how many unique pairs of RA, Dec (down to a precision) existed. This process was repeated for each pair of CDS table name and photometric band, and the maximum number of unique sources inside the searching radius was stored as representative of the neighbourhood of that source. The same procedure was repeated using a 2” searching radius as well. The distribution of the number of neighbours for sources in our catalogue is shown in Fig. 9.
This procedure allowed us to leverage the large amount of photometric data available within CDS to investigate the neighbourhood of sources in our catalogue, including data observed by instrumental setups yielding a wide variety of pixel scales and PSF sizes, while being agnostic to the selection window introduced by the criteria for inclusion in this compilation discussed in Sec. 3. We found that 6,912 sources in our catalogue have at least one neighbour within 2” from sources’ adopted coordinate, and 12,060 have neighbours within 5”. Sources with at least one neighbour within 2” were labelled as visual pair candidates (V?) with the support that 71% of these sources were also identified as such by other multiplicity labels considered.
@@ 4.5 Possible Duplicates
After extensively evaluating the multiplicity of sources in our catalogue, we are now equipped to address the incidence of duplicated sources in our data compilation process. An internal match with sources’ sky positions and a 2” radius reveals that 1,864 sources are possibly duplicated. We verified that 1,687 of these had a multiplicity flag derived in this section or had their possible duplication flagged as multiple. From the remaining, 119 sources could be traced back to the publications introducing them to our dataset in Sec. 2.1, which were typically surveys including higher-resolution images, and we could confirm that multiple sources were indeed reported by their original publication, 4 had support for being independent sources from a match with Simbad, and 20 had a neighbourhood flag (Sec. 4.4) indicative of at least one neighbour within a 5” radius. Finally, only 34 sources could not be explained and are likely duplicates due to some coordinate mismatch in our historical data compilation. These sources were flagged accordingly.
5 Further discussion
@@ 5.1 Comparison with Simbad
We used Simbad’s TAP service (on 13th December 2024) to retrieve data for 165,831 sources indexed by Simbad and located in the same field of our survey. We queried this sample using a 2” searching radius and found 28,384 Simbad sources with a possible counterpart in our catalogue. All possible counterparts were recorded and are reported in our catalogue regardless of their Simbad object type label (Tab. ).
@@ 5.1.1 Young sources known as such by Simbad
We followed SIMBAD User’s Guide (v1.8) guidelines on object types to identify 18,857 sources in the OSFC field with object_type labels indicative of T Tauri Stars (TT? or TT*), Young Stellar Object (Y*?, Y*O), Ae/Be Herbig stars (Ae* or Ae?), or Orion Variables (Or*, which variable stars with irregular variability associated to eruption phenomena in young stars). Note that this list does not include objects that are arguably related to star-forming regions such as dense cores (cor) and Herbig Haro objects (HH). Although some of these sources are indirectly included in our catalogue, as discussed in Sec. 2.1, they were considered outside the main scope of our compilation. Here, we call the first group of Simbad’s object type labels ”YSO labels”, i.e. labels that directly imply the source is a YSO. We call the second group ”star-formation labels”, i.e. labels that imply that the source is part of a star-forming region. The second set of labels adds 6,839 sources to the list of sources related to a star-forming region known by Simbad. Considering both sets of labels, 17,535 sources in our catalogue were already known by Simbad as young. Among the 7,995 sources in our catalogue with a close Sibmad counterpart but not in the two categories of labels considered, 85% were generically labelled as star (*).
@@ 5.1.2 YSOs known by Simbad but missed by our compilation
Among the YSOs known by Simbad, 2,363 were not included in our catalogue. 280 of these are located in Mon R2, which was excluded from our survey. We have retrieved the bibliographic references associated with the remaining sources and verified that 1,005 of them originate from studies with titles and abstracts focused on clumps and cores, which were considered out of our scope in Sec. 2.1. 251 sources could be traced back to CDS tables included in our compilation but were not deemed YSO candidates by the criteria in Sec. 3. The remaining sources were associated with one of the following cases. They came from all-sky variability surveys including ML-generated YSO variability flags (e.g. Samus’ et al., 2017); they came from all-sky cluster membership studies with Gaia data, whereas we have only included such surveys in our compilation when they were specifically tailored for studying the OSFC or star-forming regions (e.g. Zari et al., 2018); they came from studies focused on the OSFC but did not include tables on CDS and had too few sources for their tables to make it to the list of tables digitalised by us in Sec. 2.1.2; They came from studies focussed on YSOs, but with bibliographic entries absent of the searching terms utilised in Sec. 2.1.1. Finally, we identified one bibliographic entry that adds a few dozen YSO candidates (Muench et al., 2002), which was discarded in Sec. 2.1.2 due to its focus on massive stars.
@@ 5.1.3 Sources unknown to Simbad
Our catalogue included 2,349 sources without a counterpart in Simbad (1,977 if a 5” search radius is used). It is interesting to note that although these sources were unknown to Simbad, 81% of them had data collected with the VizieR-SED tool within a search radius of 2”. 90% of these sources had at least 4 data points in their compiled SEDs, but some had as many as 59 data points. Beyond their SEDs, 2,052 sources unknown to Simbad had measurements of at least one of the relevant observables discussed in Sec. 3, with 78% having in Sec. 3.3.1 comprising 506, 868, 157, 58, and 254 were classified as diskless Class III, thin-disk Class III, Class II, Flat-Spectrum and Class 0/I respectively. The 297 sources unknown to Simbad that did not include any relevant observables related to youth other than photometry could be traced back to publications included in our compilation that selected YSOs using a CMD, with 141 of these sources from Da Rio et al. (2009) and 77 from Suárez et al. (2019). Interestingly, 1,546 of these sources are included in Gaia DR3 and 152 of them are classified as YSOs by the Gaia DR3 variability pipeline (Marton et al., 2023)
@@ 5.2 Non-YSOs and catalogue purity
In Sec. 3, we employed a positive evidence rule to compose our catalogue, where a single peer-reviewed study reporting a source as a YSO candidate was enough to grant its inclusion here. We have thus prioritised the completeness of our YSO census, rather than its purity. Consequently, YSO purity must be addressed depending on the use intended for the catalogue. Although a formal evaluation of the reliability of youth indicators of each individual source in the catalogue is beyond the scope of this paper and the subject of a follow-up study (Roquette et al., in prep.), here we discuss a series of assumptions to assess our catalogue’s degree of contamination by different types of sources. To support this discussion, Fig. 10 shows a Gaia DR3 CMD for 23,716 sources in our sample with Gaia photometry and parallaxes. As a reference, we used solar-metallicity MESA isochrones (Dotter, 2016).
@@ 5.2.1 Massive stars
We call massive stars high-mass stars that are past the PMS phase. From a combination of the adopted MESA stellar evolution models with Pecaut & Mamajek (2013) -Spectral Type scales, and considering sources at 1 Myrs, this is equivalent to sources with spectral type earlier than B3. Based on the spectral types collated in Sec. 3.1.1, 80 sources in our catalogue met this criterion (see Fig. 10 left panel), yielding a incidence of massive stars.
@@ 5.2.2 Contamination by Giants
Because background reddened giants are a known major pollutant to YSO identification studies, we wish to further add value to our catalogue by ranking possible giant contaminants. In Sec. 3.5 we collected 57,649 measurements for 12,313 sources in our catalogue. Although giant contamination in our catalogue was already hinted at in Fig. 6, the direct use of these measurements for the purpose of identifying giants is hampered by the wide spread in derivations from different surveys. Thus, we used uncertainties as a statistical weight when combining multiple literature values and restricted our analysis to 6,895 sources with either a 3 dispersion or uncertainty (if only one measurement is available) below 0.5 dex. Fig. 10 (right panel) shows the distribution of for this sample, which yields the identification of 457 likely giants with (this threshold was chosen based on distributions of values in the MESA evolutionary models with the aim of minimising overlap in distributions of giants and YSO).
To expand the giant contamination investigation to a larger portion of our catalogue, in App. D we employed giants identified based on estimation from large spectroscopic surveys to implement a Random-Forest binary classifier trained on features from Gaia DR3 to estimate the probability that sources in our catalogue have properties comparable to a known sample of Giant stars. The results of this giant classifier are shown in the middle panel of Fig. 10. We were able to derive such probabilities for 20,714 sources in our catalogue (reported in Tab. ). We found 2,286 sources with probabilities larger than 90%, suggesting a giant contamination level of in our catalogue.
@@ 5.2.3 Extragalactic contamination
In App. D.2.1, we have collated a sample of 32,850 known extragalactic sources in the field of the OSFC including resolved and unresolved Galaxies, AGNs and QSOs. Among this literature extragalactic sample, 154 sources had a close counterpart in our catalogue. This labelled sample was used in the implementation of a Random Forest binary classifier trained to identify extragalactic sources based on photometric colours from the PanSTARRs DR2 and CatWISE surveys. The trained classifier was used to estimate the probability that sources in our catalogue have properties comparable to a known sample of extragalactic sources and allowed us to identify 110 likely extragalactic sources in our catalogue, 51 of those previously unlabelled extragalactic. Fig. 11 shows one of the parameter spaces defined by the features used in our classifier.
@@
Altogether, we identified 205 probable extragalactic sources in our sample. This accounts for expected level of contamination by extragalactic sources in our catalogue.

### v2

@@ 4.2 Multiplicity labels in from all-sky catalogues
4.3 Big data identification of visual-pair candidates
In Sect. 2.4, we have employed the VizieR-SED tool to retrieve the photometric data available on CDS around the position of each source in our catalogue. In this section, we further employ these data to gauge sources’ neighbourhoods.
Starting from the raw VizieR-SED output for a 5” searching radius, rather than relying again on the procedure described in Sect. 3.2 we carried a simpler preprocessing. We kept only data in the wavelength range 0.01-5, excluding multi-epoch surveys (e.g., TESS) and surveys with source detection and extraction at different apertures (e.g., SDSS). We then grouped the remaining data by unique CDS table name and photometric band and counted how many unique pairs of RA, Dec (down to a precision) existed. This process was repeated for each pair of CDS table name and photometric band, and the maximum number of unique sources inside the searching radius was stored as representative of the neighbourhood of that source. The same procedure was repeated using a 2” searching radius as well. The distribution of the number of neighbours for sources in our catalogue is shown in Fig. 10.
This procedure allowed us to leverage the large amount of photometric data available within CDS to investigate the neighbourhood of sources in our catalogue, including data observed by instrumental setups yielding a wide variety of pixel scales and PSF sizes, while being agnostic to the selection window introduced by the criteria for inclusion in this compilation discussed in Sect. 3. We found that 6,912 sources in our catalogue have at least one neighbour within 2” from the sources’ adopted coordinates, and 12,060 have neighbours within 5”. Sources with at least one neighbour within 2” were labelled as visual pair candidates (V?) with the support that 71% of these sources were also identified as such by other multiplicity labels considered.
@@ 4.4.3 Gaia’s Renormalised Unit Weight Error
4.4.4 Gaia’s RUWE in the context of a large sample of YSOs
As an illustrative application of our catalogue, we further contribute to probing previous suggestions that Gaia’s RUWE may be inflated in YSOs. As discussed in Sect. 4.4.3, Gaia’s astrometric solution assumes that sources are point-like with a well-defined centre of light and a single-star movement. Along with multiple systems, extended sources also deviate from these assumptions due to their non-point-like nature and the underlying larger uncertainty in resolving their centre of light. In YSOs, the diffuse emission of the envelope in Class 0/I sources, or even the radiation re-emitted by a thick disc in Class II sources could be the culprit of an uncertain centre of light.
Does the presence of a disc inflate Gaia’s RUWE?
The inflation of RUWE values due to the presence of a circumstellar disc has previously been suggested by Fitton et al. (2022), based on a modest sample of 122 single stars selected from adaptative optics surveys for which they find an excess of disc-bearing sources in the RUWE range 1–2.5. Here, we contribute to probing this evidence by examining a much larger sample of YSOs. For this end, we selected sources from our catalogue with both derived in Sect. 3.3.1, and Gaia’s DR3 RUWE.
We defined two subsamples based on (Sect. 3.3.1), namely sources with a thick disc () and discless sources (). To maximise the reliability of sample selections using the indices, we considered only sources with at least one datapoint available for m. We note that the requirement for Gaia data will inevitably remove the least evolved and most embedded Class I/O sources from our catalogue, as they are too faint and beyond Gaia’s detection capacity. To maximise sample purity, we also removed all sources flagged as possible contaminants in Sect. 5.2. Next, we also removed all sources flagged as multiple in Sect. 4, except for cases where the only multiplicity flag was the RUWE-based one, R (Sect. 4.4.3). This is required to preserve sources in the RUWE range that Fitton et al. (2022) claim to be populated by disc-bearing sources. Although 51% of the R labelled sample is still removed, as they were also identified as multiple systems by other methods, a modest fraction of sources with RUWE and probably real unresolved systems are left in the sample. To further minimise the influence of remaining binaries in our results, we focused our analysis on the RUWE2.5 range, which has been suggested to be the relevant range for disc-bearing sources (see Fitton et al., 2022). Our final sample was composed of 4 364 discless sources and 1 489 thick disc sources.
We employed a series of non-parametric statistical tests (Kolmogorov-Smirnov, Mann-Whitney U, and Anderson-Darling) to examine the differences between RUWE distributions of disc-bearing and discless sources, finding no support for significant differences between the two distributions. Alternatively, we have also explored other criteria for sample selection (e.g., the use of colours for thick disc selection) and other statistical approaches, such as examining disc fraction as a function of RUWE, but found no statistically significant indication of RUWE inflation due to the presence of thick discs around less evolved YSOs in our sample.
Does the variability of YSOs inflate RUWE?
Beyond extended sources, photometric variability has also been proposed as a contributor to RUWE. For example, Belokurov et al. (2020) previously uncovered a trend of inflated RUWE as a function of variability amplitude for RRLyrae and Cepheids observed with Gaia DR2. As further discussed by these authors, the inflation of RUWE by variability can be traced back to RUWE’s definition (Lindegren, 2018, eq. (4)). As its name says, RUWE is a normalised version of Gaia’s 5-parameter astrometric solution’s unit error, where a normalisation coefficient is derived from one of the modes UWE’s as a function of magnitude and colour. In variable stars where sharp changes of magnitude and colour take place, this normalisation may be incorrect, yielding spurious inflation of RUWE. Since variability is a prevalent characteristic of YSOs (Sect. 3.8), here we also address the influence of variability on Gaia’s DR3 RUWE.
We used Gaia DR3 calculated in Sect. 3.8.1 (available for of the sources in our catalogue) to address the evidence for variability influence on RUWE (e.g. Belokurov et al., 2020; Barber & Mann, 2023). As in the previous section, we minimised the incidence of binaries in the sample by removing all sources labelled as such, except for sources with only a R flag, and examining RUWE values up to 2.5. We also removed all sources flagged as possible contaminants in Sect. 5.2, which left us with 10,326 sources. Next, we bin the data at magnitude ranges (848 sources), (3,571), (3,850), and : (2,057) with the first two bins following Gaia’s window classes, and the remaining aimed at minimising the effects of the signal-to-noise variations with magnitude. For each magnitude bin, we estimate the first, second, and third rolling quartiles for the distributions of RUWE as functions of . The results are shown in Fig. 11, where variations of in RUWE are verified as a function of amplitude.
@@ 4.5 Possible Duplicates
After extensively evaluating the multiplicity of sources in our catalogue, we are now equipped to address the incidence of duplicated sources in our data compilation process. An internal match with sources’ sky positions and a 2” radius reveals that 1,864 sources are possibly duplicated. We verified that 1,687 of these had a multiplicity flag derived in this section or had their possible duplication flagged as multiple. From the remaining, 119 sources could be traced back to the publications introducing them to our dataset in Sect. 2.1, which were typically surveys including higher-resolution images, and we could confirm that multiple sources were indeed reported by their original publication, 4 had support for being independent sources from a match with Simbad, and 20 had a neighbourhood flag (Sect. 4.3) indicative of at least one neighbour within a 5” radius. Finally, only 34 sources could not be explained and are likely duplicates due to some coordinate mismatch in our historical data compilation. These sources were flagged accordingly.
5 Catalogue purity and contamination assessment
@@ 5.1 Comparison with Simbad
We used Simbad’s TAP service (on 13th December 2024) to retrieve data for 165 831 sources indexed by Simbad and located in the same field as our survey. We used a 5” searching radius to search for possible counterparts at the position of each source in our catalogue. This relatively large searching radius was chosen to account for possible offsets between coordinates adopted in the historical compilation and the ones adopted by CDS888It was noted during the data integration process (Sect. 2.1.3) that infrared source coordinates (e.g., from 2MASS) sometimes show offsets with optical coordinates (e.g., from Gaia) that cannot be explained by the different epochs of observations. This was particularly relevant for YSOs with large envelope contributions or part of multiple systems. As a cautionary example, we point to the star V 1118 Ori (NEMESIS Internal_ID 5465), a widely studied Class II YSO reported in our database with its 2MASS coordinate (2MASS J05344474-0533421). V 1118 Ori is indexed by Simbad with its ICSR J2000 coordinates from Gaia DR2, which are offset by 3.67” from its 2MASS coordinates.. We found 33 846 sources indexed by Simbad with a possible counterpart in our catalogue. 25 902 sources in our catalogue have at least one Simbad counterpart within the searching radius. All possible counterparts were recorded and are reported in our catalogue, regardless of their Simbad object type label (Tab. ), including 3 167 sources with multiple possible counterparts. We leave it to the users of the catalogue to rank possible counterparts according to their use case.
@@ 5.1.1 Young sources known as such by Simbad
We followed SIMBAD User’s Guide (v1.8) guidelines on object types to identify 19 625 sources in the OSFC field with object_type labels indicative of T Tauri Stars (TT? or TT*), Young Stellar Object (Y*?, Y*O), Ae/Be Herbig stars (Ae* or Ae?), or Orion Variables (Or*, which variable stars with irregular variability associated with eruption phenomena in young stars). We note that this list does not include objects that are arguably related to star-forming regions, such as dense cores (cor) and Herbig Haro objects (HH). Although some of these sources are indirectly included in our catalogue, as discussed in Sect. 2.1, they were considered outside the main scope of our compilation. Here, we call the first group of Simbad’s object type labels ”YSO labels”, i.e. labels that directly imply the source is a YSO. We call the second group ”star-formation labels”, i.e. labels that imply that the source is part of a star-forming region. Considering both sets of labels, 18 029 sources in our catalogue were already known by Simbad as young. Among the 7 873 sources in our catalogue with a close Sibmad counterpart but not in the two categories of labels considered, 87% were generically labelled as star (*).
@@ 5.1.2 YSOs known by Simbad but missed by our compilation
Among the YSOs known by Simbad, 2 112 were not included in our catalogue. 280 of these are located in Mon R2, which was excluded from our survey. We have retrieved the bibliographic references associated with the remaining sources and verified that 1,005 of them originate from studies with titles and abstracts focused on clumps and cores, which were considered out of our scope in Sect. 2.1. 251 sources could be traced back to CDS tables included in our compilation but were not deemed YSO candidates by the criteria in Sect. 3. The remaining sources were associated with one of the following cases. They came from all-sky variability surveys including ML-generated YSO variability flags (e.g. Samus’ et al., 2017); they came from all-sky cluster membership studies with Gaia data, whereas we have only included such surveys in our compilation when they were specifically tailored for studying the OSFC or star-forming regions (e.g. Zari et al., 2018); they came from studies focused on the OSFC but did not include tables on CDS and had too few sources for their tables to make it to the list of tables digitalised by us in Sect. 2.1.2; They came from studies focused on YSOs, but with bibliographic entries absent from the searching terms utilised in Sect. 2.1.1. Finally, we identified one bibliographic entry that adds a few dozen YSO candidates (Muench et al., 2002), which was discarded in Sect. 2.1.2 due to its focus on massive stars.
@@ 5.1.3 Sources unknown to Simbad
Our catalogue included 1 977 sources without a counterpart in Simbad. It is interesting to note that although these sources were unknown to Simbad, 90% of them had data collected with the VizieR-SED tool within a search radius of 2”. 95% of these sources had at least 4 data points in their compiled SEDs, but some had as many as 50 data points. Beyond their SEDs, 1 623 sources unknown to Simbad had measurements of at least one of the relevant observables discussed in Sect. 3, with 89% having in Sect. 3.3.1 comprising 477, 847, 140, 55, and 237 were classified as discless Class III, thin-disc Class III, Class II, Flat-Spectrum and Class 0/I respectively. The 125 sources unknown to Simbad that did not include any relevant observables related to youth other than photometry could be traced back to publications included in our compilation that selected YSOs using a CMD, with 42 of these sources from Da Rio et al. (2009), 70 from Suárez et al. (2019) and 42 from Bouy et al. (2014).
@@ 5.2 Non-YSOs and catalogue purity
In Sect. 3, we employed a positive evidence rule to compose our catalogue, where a single peer-reviewed study reporting a source as a YSO candidate was enough to grant its inclusion here. We have thus prioritised the completeness of our YSO census, rather than its purity. Consequently, YSO purity must be addressed depending on the intended use of the catalogue. Although a formal evaluation of the reliability of youth indicators of each individual source in the catalogue is beyond the scope of this paper and the subject of a follow-up study, here we discuss a series of assumptions to assess our catalogue’s degree of contamination by different types of sources. To support this discussion, Fig. 12 shows a Gaia DR3 CMD for 23,716 sources in our sample with Gaia photometry and parallaxes. As a reference, we used solar-metallicity MESA isochrones (Dotter, 2016).
@@ 5.2.1 Massive stars
We call massive stars high-mass stars that are past the PMS phase. From a combination of the adopted MESA stellar evolution models with Pecaut & Mamajek (2013) -Spectral Type scales, and considering sources at 1 Myrs, this is equivalent to sources with spectral type earlier than B3. Based on the spectral types collated in Sect. 3.1.1, 80 sources in our catalogue met this criterion (see Fig. 12 left panel), yielding a incidence of massive stars.
@@ 5.2.2 Contamination by Giants
Because background reddened giants are a known major pollutant to YSO identification studies, we wish to further add value to our catalogue by ranking possible giant contaminants. In Sect. 3.5 we collected 57 649 measurements for 12 313 sources in our catalogue. Although giant contamination in our catalogue was already hinted at in Fig. 7, the direct use of these measurements for the purpose of identifying giants is hampered by the wide spread in derivations from different surveys. Thus, we used uncertainties as a statistical weight when combining multiple literature values and restricted our analysis to 6 895 sources with either a 3 dispersion or uncertainty (if only one measurement is available) below 0.1 dex. Fig. 12 (right panel) shows the distribution of for this sample, which yields the identification of 537 likely giants with (this threshold was chosen based on distributions of values in the MESA evolutionary models with the aim of minimising overlap in distributions of giants and YSO).
To expand the giant contamination investigation to a larger portion of our catalogue, in App. E we employed giants identified based on estimation from large spectroscopic surveys to implement a Random-Forest binary classifier trained on features from Gaia DR3 to estimate the probability that sources in our catalogue have properties comparable to a known sample of Giant stars. The results of this giant classifier for our sample are shown in the middle panel of Fig. 12. We were able to derive such probabilities for 20 546 sources in our catalogue (reported in Tab. ), and 2 614 sources with high probabilities of being a Giant. We thus estimate a giant contamination level of in our catalogue. We note that among high giant probability sources, 515 sources were in the sample detected from their logg. Altogether, sources with logg indicative of giants and sources identified by our classifier correspond to 9.5% of our catalogue, suggesting a small residual incidence of unlabelled giants.
@@ 5.2.3 Extragalactic contamination
In App. E.2.1, we collated a sample of known extragalactic sources in the field of the OSFC, including resolved and unresolved Galaxies, AGNs and Quasi-stellar objects or Quasars (QSOs). Among this literature-based extragalactic sample, 34 sources had a close counterpart in our catalogue. The full labelled sample was used in the implementation of a Random Forest binary classifier trained to identify extragalactic sources based on photometric colours from the Gaia DR3, PanSTARRs DR2, and CatWISE surveys. The trained classifier was used to evaluate the probability of 15 855 sources in our catalogue having properties comparable to the known sample of extragalactic sources. This allowed us to identify 140 likely extragalactic sources in our catalogue, 115 of those previously unlabelled extragalactic, yielding an estimated incidence of extragalactic contaminant in our sample. Fig. 13 shows one of the parameter spaces defined by the features used in our classifier. Altogether, we identified 159 probable extragalactic sources in our sample (, suggesting a small residual incidence of extragalactic contaminants.
We note that the misclassification of extragalactic sources and YSOs happens both ways. During the collation of a labelled sample for our extragalactic classifier in App. E, we refrained from including the Gaia DR3 QSO candidates in the labelled sample due to evidence of a relevant incidence of YSOs in this sample. Following Gaia DR3’s documentation (Gaia Collaboration et al., 2023b) for selecting a high-purity QSO sample towards the OSFC, we found 81 QSO candidates in common with our YSO catalogue. Further examination of these sources on light of the diverse attributes included in our database revealed that a relevant fraction of such sources had compelling evidence for bona fide YSO classification in our catalogue, with 35 of these having spectroscopic constraints indicative of youth. Out of the 46 remaining, 30 were identified as high probability extragalactic contaminants in App. E, but 16 are unaccounted. While we opt not to include this sample in our contaminant sample, users of the catalogue who wish for a full removal of possible extragalactic sources are recommended to use the extragalactic_label and strat_label (Tab. ) to identify and further inspect such sources.